OpenStudy (anonymous):

in circular motion why is the formula for velocity 2πr/T i dont understand what multiplying it by the radius does that angular velocity does not already answer....is it then m/s rather than radians?

OpenStudy (anonymous):

2πr=S - the circumference or way, t - time, then for usualy velocity v=S/t=2πr/t